OpenStudy (shamil98):

Double integral: \[\huge \int\limits_{}^{}\int\limits_{R}^{} 2x-4y^3~dA, R = [-5,4] \times [0,3]\]

terenzreignz (terenzreignz):

Well... this doesn't look too bad... x goes from -5 to 4, y goes from 0 to 3 am I right?

terenzreignz (terenzreignz):

Oi @shamil98 work with me here XD \[\Large \iint\limits_R2x-4y^3 dA\]

OpenStudy (shamil98):

Yeah. what would be the first step tho?

terenzreignz (terenzreignz):

Here... now that we have the limits down, the rest is easy (sort of) \[\Large \int\limits_{y=0}^3 \ \int\limits_{x=-5}^42x-4y^3 dxdy\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (shamil98):

Yeah.

terenzreignz (terenzreignz):

Now, much like with partial differentiation, you integrate this part first \[\Large \int\limits_{y=0}^3 \ \color{blue}{\int\limits_{x=-5}^42x-4y^3 dx}dy\] While for the time being, treating y as a constant.

OpenStudy (shamil98):

A bit lost, what would you integrate -4y^3 as ? ...

terenzreignz (terenzreignz):

That entire thingy is just constant (with respect to x) so it's simply \(\large -4xy^3\)

OpenStudy (shamil98):

Oh

terenzreignz (terenzreignz):

that was your problem, I assume? :D

OpenStudy (luigi0210):

*following along*

OpenStudy (shamil98):

yeah, i was wondering that lol

terenzreignz (terenzreignz):

So... got it from here?

OpenStudy (shamil98):

Next , you integrate with respect to y ?.. or do you evaluate it as a definite integral ? just started this topic

terenzreignz (terenzreignz):

There are limits... so... definite integral? LOL

terenzreignz (terenzreignz):

\[\Large \int\limits_{y=0}^3 \ \color{blue}{\int\limits_{x=-5}^42x-4y^3 dx}dy\]

OpenStudy (primeralph):